OpenStudy (anonymous):

Graph y=3(x+2)^2-4 Label all parts on the graph and 5 points. Please help!!

OpenStudy (anonymous):

Creating a Plot The plot function has different forms, depending on the input arguments. If y is a vector, plot(y) produces a piecewise linear graph of the elements of y versus the index of the elements of y. If you specify two vectors as arguments, plot(x,y) produces a graph of y versus x. For example, these statements use the colon operator to create a vector of x values ranging from 0 to 2π, compute the sine of these values, and plot the result: x = 0:pi/100:2*pi; y = sin(x); plot(x,y) Add axis labels and a title. xlabel('x = 0:2\pi') ylabel('Sine of x') title('Plot of the Sine Function','FontSize',12) The characters \pi create the symbol π and the FontSize property increases the size the text used for the title: